OpenStudy (butterflyhope):

Urgent *will medal*: Use the following info for problems 1 and 2: The sum of the measures of the angles in any triangle is 180. If the largest angle is twice the size of the smallest and the middle angle is 1.5 times the size of the smallest, then what are the measures of each angle? 1. Let x represent the measure of the smallest angle. Write an equation to solve the problem. 2. Find the measure of each angle. Thanks c:

OpenStudy (anonymous):

is \(x\) is the smallest angle, then the middle one is \(1.5x\) and the larger one is \(2x\)

OpenStudy (anonymous):

add \[x+1.5x+2x=4.5x\]

OpenStudy (anonymous):

set that equal to \(180\) and solve for \(x\) in one step (divide by \(4.5\))

OpenStudy (anonymous):

1) x + 1.5x + 2x = 180. 2) x + 1.5x + 2x = 180 4.5x = 180 x = 180/4.5 x = 40 degree For smallest angle, the answer is 40 degree. For medium angle, the answer is 1.5(40). So, the answer is 60 degree. For large angle, the answer is 2(40). So the answer is 80 degree.

OpenStudy (anonymous):

btw not loving the new doctor does it get better?

OpenStudy (butterflyhope):

@satellite73, first thanks, and I think The Doctor will get better, I actually really like him, but he has his moments of stupidity xD Dark Water was my fave episode, although it should be renamed "Tearing Everyone's Heart Out". MOFFATTT. xD @Mirahalfian, thanks :)